OpenStudy (anonymous):

Please help!! how can I simplify 5 1/3+(-3 9/18) ? I also need to show my work and I have no idea how

OpenStudy (anonymous):

Could you write that in a clearer form? I'm not sure if you mean 51/3 or 5.33333, etc.

OpenStudy (anonymous):



OpenStudy (anonymous):

the 1/3 is a fraction and also 9/18 is a fraction

OpenStudy (anonymous):

\[5\frac{ 1 }{ 3 } + \left( -3\frac{ 9 }{18 } \right)\] So this?

OpenStudy (anonymous):

yes I just didn't know how to draw it on here @williamnz

OpenStudy (anonymous):

Ok we will first turn it into improper fraction form

OpenStudy (anonymous):

so how would I do that? @williamnz

OpenStudy (wolf1728):

The lowest common denominator is 6 and so the equation becomes 5 (2/6) -3 (3/6) Can you solve it from there?

OpenStudy (anonymous):

For the numerator, Times the 5 by the denominator, add 1 and divide by the demoninator So for the first one it would be 16/3 Can you find the second one?

OpenStudy (anonymous):

im so confused um okay so I multiply the 5 and the 3 and itll be 5 then I add 1 but why would I add one to it aswell @williamnz

OpenStudy (anonymous):

@wolf1728 how did you end up getting that equation?

OpenStudy (anonymous):

@fly_momma97 That's just the way we find improper fractions. Wolf's way is actually quicker and easier so you should stick to his

OpenStudy (wolf1728):

I looked at the denominators and found that they both could be converted to 6.

OpenStudy (anonymous):

what do you mean by convert? like multiplied?

OpenStudy (wolf1728):

okay let's just work with fractions. Let's suppose the problem is 1/3 - 9/18 How would you go about solving that?

OpenStudy (anonymous):